# Unified Theory of Magnetization Temperature Dependence in Ferrimagnetics

### v1

@@ III.2 Spin-wave supression of the YIG magnetization
$\begin{split}\overline{S}^{z}=\,&\Big{[}\sum_{j=1}^{n_{+}}\big{(}S-\langle a_{jn}^{\dagger}a_{jn}\rangle\big{)}\\ &-\sum_{j=n_{+}+1}^{n}\big{(}S-\langle a_{jn}^{\dagger}a_{jn}\rangle\big{)}\Big{]}\ .\end{split}$ (20a)
@@
$\begin{split}\overline{S}^{z}(T)=\overline{S}^{z}(0)-\sum_{\bm{k}}\Big{[}&\sum_{j=1}^{n_{+}}\langle a_{j\bm{k}}^{\dagger}a_{j\bm{k}}\rangle-\sum_{j=n_{+}+1}^{n}\langle a_{j\bm{k}}^{\dagger}a_{j\bm{k}}\rangle\Big{]}\,,\end{split}$ (20b)
@@
The Eq. (21) represents a simple and significant conclusion, announced above: the excitation of one ferromagnetic magnon (with frequencies $\omega_{\rm d}(\bm{k})$ and $\omega_{\rm da}(\bm{k})$, represented by green lines in Fig. 3) decreases the mean spin of YIG by unity. In contrast, the excitation of one antiferromagnetic magnon (with frequencies $\omega_{\rm a}(\bm{k})$ and $\omega_{\rm ad}(\bm{k})$, indicated by blue lines in Fig. 3) increases the mean spin by uinity. To determine the corrections to the temperature dependence of YIG magnetization due to magnons in thermodynamic equilibrium, we need to substitute in (21) the Bose-Einstein distribution for their occupation numbers:
@@
where $\omega_{j}(\bm{k})$ are the magnon frequency spectra. These spectra have been the focus of numerous studies, starting with a paper by Douglass in 1960 [29], followed by a report by Harris in 1963 [30], among others. Cherepanov, Kolokolov, and L’vov conducted a comprehensive analytical study of the 20 magnon branches in YIG, as detailed in their 1993 publication [23]. In the graphical form, the result of Ref. [23] is shown in Fig. 3. However, the full expressions for the magnon frequencies provided in Eq. (6) of [23] are very cumbersome. For our needs, we hope that using the simplified version detailed in Appendix D will be sufficient.
@@
In Fig. 5, we plot the temperature dependence of the mean spin $\overline{S}^{z}(T)$, represented by the solid green line ⑦, calculated from Eq. (20b) (see also this line in Fig. 2). This calculation was performed using $T$-independent dispersion relations $\omega_{j}(k)$, as explained in Appendix D. Formally, Eq. (20b) is valid in the low-temperature limit, where $t\ll T_{{}_{\rm C}}\approx 560\,\text{K}$. Interestingly, the theoretical dependence closely matches the experimental results of $\overline{S}^{z}(T)$ up to about $T\simeq 400\,\text{K}$. Recently, a similar result was obtained in [31] (compare the blue line in their Fig. 2 with our green line ⑦ in Fig. 5) through direct numerical simulation of the Landau-Lifshitz equation [22] supplemented by a stochastic random force. This force is governed by the fluctuation-dissipation theorem [22], mimicking the effect of the quantum thermostat with the Bose-Einstein distribution (24).
Note that all experiments were conducted with the external magnetic field $H_{\text{ext}}$, which creates an internal magnetic field in the sphere given by $H_{\text{int}}=H_{\text{ext}}-{4\pi M}/{3}$. This field introduces an additional frequency, $\omega_{0}=\gamma H_{\text{int}}$, for FM modes, while subtracting this frequency from the AFM modes. In terms of temperature, $\omega_{0}\approx 7.4\,\text{K}$ for $H=1\,\text{T}$. This value is relatively low compared to typical $\omega_{j}$ frequencies, which are usually of the order of a few hundred Kelvin. Consequently, the external magnetic field does not significantly alter the resulting $M(T)$ dependence (by no more than four percent). In our semiquantitative approach, this difference can be safely ignored.
@@ III.3 Temperature dependence of the spin-wave spectra
All expressions for the frequencies analysed in Appendix D are derived using the Holshtein-Primakoff representation, which is valid for the temperature $T=0$. To apply them in the intermediate temperature range, especially near the Curie temperature, one must account for the temperature dependence of all 20 spin-wave spectra branches $\omega_{j}(\bm{k})$. A relatively straightforward and simple approach to solve this problem can be derived from the simplest version of the Green’s function splitting, as suggested by Tyablikov [32]. In this method, all analytical expressions for $\omega_{j}(\bm{k})$ are multiplied by the factor $\overline{S}^{z}(T)/\overline{S}^{z}(0)$, which equals unity as $T\to 0$ and vanishes when $T\to T_{{}_{\rm C}}$.
@@
This result can be understood from a physical perspective, at least for small $a\,k\ll 1$ and relatively small temperatures $T\ll T_{{}_{\rm C}}$. Under these conditions, the main contribution to the reduction of $\overline{S}^{z}(T)$ comes from the rapid spin waves with $a\,k\sim 1$. This allows us to average the problem over fast motions, and to consider slow spin waves with $a\,k\ll 1$, at least in the limit $T\to 0$.
The yellow solid line ⑪ in Fig. 5 illustrates the temperature dependence of the normalized total mean spin $\overline{S}^{z}(T)/\overline{S}^{z}(0)$, Eq. (21), using temperature dependent frequencies $\omega_{j}(k,T)$, Eq. (25), obtained with the experimental dependence of the mean normalized spin  [21] [orange dots ②]. It crosses the value $\overline{S}^{z}(T)=0$ at $T_{\times}\approx 410\,$K, which is below the experimental Curie temperature $T^{\rm Exp}_{{}_{\rm C}}\approx 560\,$K. In contrast, a solid green line ⑦ in Figs. 2 and 5, found with temperature-independent $\omega_{j}(k)$, yields $T_{\times}\approx 780\,$K, significantly above $T^{\rm Exp}_{{}_{\rm C}}$. Clearly, accounting for the temperature dependence of $\omega_{j}(\bm{k},T)$ is necessary. However, the simple Eq. (25) overestimates the suppression of magnetization due to the excitation of spin-waves.
@@
When $\delta=1$, Eq. (27) coincides with Eq. (25), while for $\delta=0$, $\omega_{j}(\bm{k},T)$ in Eq. (27) becomes independent of temperature. For illustration, we plotted in Fig. 5 the magenta solid line ⑫ for $\overline{S}^{z}(T)$ dependence with $\delta=0.5$.
@@ IV Unifying the spin-wave and mean-field approximations
In the simplest case of one-sublattice ferromagnetics, this program was implemented in Ref. [6] and is briefly outlined in the next Section  IV.1. We extend this program to multi-sublattice ferrimagnetics, such as YIG, in Sec.  IV.2.
@@ IV.1 Unified SW and MFA model
in simple single-sublattice ferromagnetics
The original MFA does not consider the spin-wave correction to the average spin. An effective method for improving this was discussed in detail in Ref. [6], focusing on the case of ferromagnetic materials EuO and EuS, which have a single magnetic ion in their elementary cell. In [6], Eq. (4c) in the original MFA was relpaced by the
@@
IV.2 Unified SW-MFA minimal model in multi-sublattice ferrimagnetics
The goal of this section is to suggest the simplest model for unifying MFA with SWA in multi-sublattice collinear ferrimagnetics, such as YIG, while still accounting for the fundamental physics involved.
@@ IV.2 Unified SW-MFA minimal model in multi-sublattice ferrimagnetics
This is the exact relation given by (20b) for the spin-wave correction to the mean spin in YIG, which was not included in the original MFA. As discussed earlier, the dependence of $\overline{S}^{z}(T)$, shown in Fig. 5 by solid green line ⑦, aligns well with experimental observations in YIG from $T=0$ up to approximately $400\,K$.
Our next objective is to derive a new expression for the Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF-SW}}$ in multi-sublattice ferrimagnetics, such as YIG. At this temperature, the average spin $\overline{S}^{z}(T)$ approaches zero as the temperature $T$ approaches the Curie temperature from below. Note that as $\overline{S}^{z}\to 0$, and, consequently, according to (25), $\omega_{j}(\bm{k},T)\to 0$, the occupation numbers $\langle n_{j}(k)\rangle\to\infty$. In the considered case, the Bose-Einstein distribution (24) can be approximated by the Rayleigh-Jeans distribution (26) in which we have to account for temperature dependence of $\omega(k,T)$ according to Eq. (27):
@@
Recall, for YIG, $n_{+}=12$, $n_{-}=8$ and $\Delta n\equiv n_{+}-n_{-}=4$. Using the magnon frequencies from [25, 26] (cf. the right part of Tab. 2), we fix the tuning parameter $\delta$ in (33) such that $T_{{}_{\rm C}}^{{}^{\rm MF-SW}}$ agrees with the experimental value of the Curie temperature. For $\delta=0.428$, $T_{{}_{\rm C}}^{{}^{\rm MF-SW}}\approx 560\,$K in full agreement with the experiment. Recall that $\delta=0$ in (27) corresponds to the temperature-independent $\omega_{j}(\bm{k})$, while for $\delta=1$ $\omega_{j}(\bm{k},T)$ is strongly dependent on $T$ according to (25).
To determine the temperature dependence of $\overline{S}^{z}(T)$ in the unified SW-MFA minimal model (30), we numerically solved this equation using ${\cal N_{\rm d}}$ and ${\cal N_{\rm a}}$ from Eqs. (21) and (24), along with temperature-dependent frequencies from Eq. (27) with $\delta=0.428$. For more details on the solution procedure, see Appendix F. The result, shown in Fig. 2 as the red solid line ⑧, was compared to the experimental data represented by ①, ②, and ③. Taking into account that the simple model Eq. (30) has only one tuning parameter $\delta$, we find that the quantitative agreement between the theory and experiment across the entire temperature range from $T=0$ up to the Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF-SW}}$ is more than satisfactory. It suggests that the minimal model (30) adequately captures the basic physics of the problem.
@@ V Discussion and Summary
We generalized both approaches to the multi-sublattice collinear ferromagnetic. In the MFA framework, we accounted for two magnetic sub-lattices and derived coupled equations Eq. (15) for the $T$- dependence of the mean spins for each sublattice, and the analytical expression in Eq. (16a) for the Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF}}$.
Within SWA, we accounted for multiple magnon frequency spectra branches and derived Eq. (20b), which describes the suppression of $\overline{S}^{z}(T)$ due to spin-wave excitations. Given that the exact derivation of the temperature dependence of the magnon frequencies via the diagrammatic technique is too cumbersome for the level of the approximations that we used in this work, we suggested a phenomenological description of $\omega_{j}(\bm{k},T)$, Eq. (27), that employs one tuning parameter $\delta$.
The central theoretical part of our paper focuses on the unifying SW and MF approximations for the multi-sublattice ferrimagnetics. This approach culminates in Eq. (30) which connects $\overline{S}^{z}(T)$ with the total amount of magnons in the ferromagnetic and antiferromagnetic branches. Formally speaking, Eq. (30) is a relatively simple generalization of unified SW and MFA Eq. (28a) for the simple one-sublattice ferromagnetics such as EuO or EuS analysed in Ref. [6]. In Eq. (30), we replaced the spin of a single ion $S$ in EuO with the total spin $\overline{S}^{z}(0)$ in the elementary cell of a multi-sublattice ferrimagnetic. Additionally, we substituted the total number of magnons $\mathcal{N}$ in a single FM branch with the value $\mathcal{N}_{+}-\mathcal{N}_{-}$, which accounts for magnons from all FM and AFM branches in the multi-sublattice ferrimagnetics. These replacements are not justified by some consistent procedure that is expected to lead to much more complicated basic equations. A price for simplicity of the current version of a theory for magnetization in multi-sublattice ferrimagnetics is the tuning parameter $\delta$ in Eq. (27) for $\omega_{j}(\bm{k},T)$, which was chosen to require the theoretical Curie temperature to match its experimental value.
The unified SW-MFA minimal model with the chosen tuning parameter $\delta=0.428$ demonstrate good quantitative agreement between the theoretical $\overline{S}^{z}(T)$ and experimental magnetization temperature dependences of YIG in the entire temperature range from $T=0$ K to $T=T^{\rm Exp}_{\rm{}_{C}}$.
@@
To the best of our knowledge, this is the first demonstration that $M(T)$ scales in YIG nearly normally as $\sqrt{T_{\rm{}_{C}}-T}$ for all temperatures. This behavior contrasts sharply with the anomalous scaling of $M(T)$ in simple ferromagnets EuO and EuS, where $\beta\approx 1/3$[6]. This fact highlights the need for further investigation of the magnetization of different materials.
Our theoretical approach marks a significant advancement in the description of magnetic systems. We demonstrated the effectiveness of unifying two simple theoretical approximations for the description of complex magnetic materials. We believe that the applicability of our SW-MFA minimal model is not limited to YIG but may be extended to various multi-sublattice ferromagnetics and antiferromagnetics. A further generalization of our minimal model, incorporating dipole-dipole interactions, will allow us to study more complex magnetic structures, such as Gadolinium Gallium Garnet (GGG, Gd3Ga5O12) at low temperatures [35, 36, 37, 38]. GGG has the same crystal structure and lattice parameters as YIG, making it an ideal substrate material for the epitaxial films of YIG. These layers have recently been used in research on magnon lifetime and transport at millikelvin temperatures [39, 16, 40, 41], paving the way for advancements in quantum magnonics.

### v2

@@ III.2 Spin-wave supression of the YIG magnetization
$\begin{split}\overline{S}^{z}(T)=\,&\Big[\sum_{j=1}^{n_{+}}\big(S-\langle a_{jn}^{\dagger}a_{jn}\rangle\big)\\ &-\sum_{j=n_{+}+1}^{n}\big(S-\langle a_{jn}^{\dagger}a_{jn}\rangle\big)\Big]\ .\end{split}$ (20a)
@@
[…]
@@
The Eq. (21) represents a simple and significant conclusion, announced above: the excitation of one ferromagnetic magnon (with frequencies $\omega_{\rm d}(\bm{k})$ and $\omega_{\rm da}(\bm{k})$, represented by green lines in Fig. 3) decreases the mean spin of YIG by unity. In contrast, the excitation of one antiferromagnetic magnon (with frequencies $\omega_{\rm a}(\bm{k})$ and $\omega_{\rm ad}(\bm{k})$, indicated by blue lines in Fig. 3) increases the mean spin by unity. To determine the corrections to the magnon-induce corrections to the temperature dependence of YIG magnetization in thermodynamic equilibrium, we need to substitute in (21) the Bose-Einstein distribution for their occupation numbers:
@@
where $\omega_{j}(\bm{k})$ are the magnon frequency spectra.
These spectra have been the focus of numerous studies, starting with a paper by Douglass in 1960 [32], followed by a report by Harris in 1963 [33], among others. Cherepanov, Kolokolov, and L’vov conducted a comprehensive analytical study of the 20 magnon branches in YIG, as detailed in their 1993 publication [23]. In the graphical form, the result of Ref. [23] is shown in Fig. 3. However, the full expressions for the magnon frequencies provided in Eq. (6) of [23] are very cumbersome. For our needs, we hope that using the simplified version detailed in Appendix F will be sufficient.
@@
In Fig. 5, we plot the temperature dependence of the mean spin $\overline{S}^{z}(T)$, represented by the solid green line ⑦, calculated from Eq. (20b) (see this line also in Fig. 2). This calculation was performed using $T$-independent dispersion relations $\omega_{j}(k)$, as explained in Appendix H. Formally, Eq. (20b) is valid in the low-temperature limit, where $t\ll T_{{}_{\rm C}}\approx 560\,\text{K}$. Interestingly, the theoretical dependence closely matches the experimental results of $\overline{S}^{z}(T)$ up to about $T\simeq 400\,\text{K}$ Recently, a similar result was obtained in [34] (compare the blue line in their Fig. 2 with our green line ⑦ in Fig. 5) through direct numerical simulation of the Landau-Lifshitz equation  [22] supplemented by a stochastic random force. This force is governed by the fluctuation-dissipation theorem  [22], mimicking the effect of the quantum thermostat with the Bose-Einstein distribution  (24).
Note that all experiments were conducted with the external magnetic field $H_{\text{ext}}$ which creates an internal magnetic field in the sphere given by $H_{\text{int}}=H_{\text{ext}}-{4\pi M}/{3}$. This field introduces an additional frequency, $\omega_{0}=\gamma H_{\text{int}}$ for FM modes, while subtracting this frequency from the AFM modes. In terms of temperature, $\omega_{0}\approx 7.4\,\text{K}$for $H=1\,\text{T}$ This value is relatively low compared to typical $\omega_{j}$ frequencies, which are usually of the order of a few hundred Kelvin. Consequently, the external magnetic field does not significantly alter the resulting $M(T)$ dependence (by no more than four percent). In our semi-quantitative approach, this difference can be safely ignored.
@@ III.3 Temperature dependence of the spin-wave spectra
All expressions for the frequencies analyzed in Appendix F are derived using the Holshtein-Primakoff representation, which is valid for the temperature $T=0$ K. To apply them in the intermediate temperature range, especially near the Curie temperature, one must account for the temperature dependence of all 20 spin-wave spectra branches $\omega_{j}(\bm{k})$. A relatively straightforward and simple approach to solve this problem can be derived from the simplest version of the Green’s function splitting, as suggested by Tyablikov [35]. In this method, all analytical expressions for $\omega_{j}(\bm{k})$ are multiplied by the factor $\overline{S}^{z}(T)/\overline{S}^{z}(0)$, which equals unity as $T\to 0$ and vanishes when $T\to T_{{}_{\rm C}}$.
@@
This result can be understood from a physical perspective, at least for small $a\,k\ll 1$ and relatively small temperatures $T\ll T_{{}_{\rm C}}$. Under these conditions, the main contribution to the reduction of $\overline{S}^{z}(T)$ comes from the rapid spin waves with $a\,k\sim 1$. This allows us to average the problem over fast motions, and to consider slow spin waves with $a\,k\ll 1$, at least in the limit $T\to 0$ K. Therefore, the classical Holstein-Primakoff representation, which is valid at $T=0$ K, should be replaced by its temperature-modified version for non-zero temperatures. In this modification, $S$ is replaced by $S_{z}(T)$. When this modified version of the Holstein-Primakoff representation is directly substituted into the exchange Hamiltonian, it reveals that the quadratic terms involving $a$ and $a^{\dagger}$ are proportional to $\overline{S}^{z}(T)$ multiplied by a linear combination of the exchange integrals. This indicates that the same structure — $\overline{S}^{z}(T)$ multiplied by the exchange integrals — will govern the magnon frequencies in ferromagnets and ferrimagnets, including YIG. For instance, the expression for the frequency $\omega(k)$ in Eq.(29) below becomes $\omega(k,T)=\overline{S}^{z}(T)(J_{0}-J_{k})$. This result is consistent with Eq.(25).
The yellow solid line ⑪ in Fig. 5 illustrates the temperature dependence of the normalized total mean spin $\overline{S}^{z}(T)/\overline{S}^{z}(0)$, Eq. (21), using temperature dependent frequencies $\omega_{j}(k,T)$, Eq. (25), obtained with the experimental dependence of the mean normalized spin  [21] [orange dots ②]. It crosses the value $\overline{S}^{z}(T)=0$ at $T_{\times}\approx 410\,$K, which is below the experimental Curie temperature $T^{\rm Exp}_{{}_{\rm C}}\approx 560\,$K. In contrast, a solid green line ⑦ in Figs. 2 and 5, found with temperature-independent $\omega_{j}(k)$, yields $T_{\times}\approx 780\,$K, significantly above $T^{\rm Exp}_{{}_{\rm C}}$. Clearly, accounting for the temperature dependence of $\omega_{j}(\bm{k},T)$ is necessary. However, the simple Eq. (25) overestimates the suppression of magnetization casued by the excitation of spin-waves.
@@
[…]
We note that Eq. (27) was proposed as an “educated guess”, justified solely by the successful comparison of the resulting $(M(T)$ with experimental data. The analytical derivation of this equation (if possible), along with a direct comparison of the proposed dependence $\omega(\bm{k},T)$ with experimental observations, will be addressed in the future. Additional details about the differences in our approach to EuO and YIG can be found in Appendix B.
@@ IV Unifying the spin-wave and mean-field approximations
In the simplest case of one-sublattice ferromagnets, this program was implemented in Ref. [6] and is briefly outlined in the next Section  IV.1. We extend this program to multi-sublattice ferrimagnets, such as YIG, in Sec.  IV.2.
@@ IV.1 Unified SW and MFA model
in simple single-sublattice ferromagnets
The original MFA does not consider the spin-wave correction to the average spin. An effective method for improving this deficiency was discussed in detail in Ref. [6], focusing on the case of ferromagnetic materials EuO and EuS, which have a single magnetic ion in their elementary cell. In [6], Eq. (4c) in the original MFA was replaced by the
@@
IV.2 Unified SW-MFA minimal model in multi-sublattice ferrimagnets
The goal of this section is to suggest the simplest model for unifying MFA with SWA in multi-sublattice collinear ferrimagnets, such as YIG, while still accounting for the fundamental physics involved.
@@ IV.2 Unified SW-MFA minimal model in multi-sublattice ferrimagnets
This equation coincides with a single-sublattice version of the Unified SW and MFA model (28a) if one replaces $\Delta n$ with unity and ${\cal N}_{+}-{\cal N}_{-}$ with $\cal N$}.
@@
This is the exact relation given by (20b) for the spin-wave correction to the mean spin in YIG, which was not included in the original MFA. As discussed earlier, the dependence of $\overline{S}^{z}(T)$, shown in Fig. 5 by solid green line ⑦, aligns well with experimental observations in YIG from $T=0$ up to approximately 400K400\,K.
Our next objective is to derive a new expression for the Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF-SW}}$ in multi-sublattice ferrimagnets, such as YIG. At this temperature, the average spin $\overline{S}^{z}(T)$ approaches zero as the temperature $T$ approaches the Curie temperature from below. Note that as $\overline{S}^{z}\to 0$, and, consequently, according to (25), $\omega_{j}(\bm{k},T)\to 0$, the occupation numbers $\langle n_{j}(k)\rangle\to\infty$. In the considered case, the Bose-Einstein distribution (24) can be approximated by the Rayleigh-Jeans distribution (26) in which we have to account for temperature dependence of $\omega(k,T)$ according to Eq. (27):
@@
Recall, for YIG, $n_{+}=12$, $n_{-}=8$ and $\Delta n\equiv n_{+}-n_{-}=4$. Using the magnon frequencies from [28, 29] (cf. the right part of Tab. 2), we fix the tuning parameter $\delta$ in (33) such that $T_{{}_{\rm C}}^{{}^{\rm MF-SW}}$ agrees with the experimental value of the Curie temperature. For $\delta=0.428$, $T_{{}_{\rm C}}^{{}^{\rm MF-SW}}\approx 560\,$K in full agreement with the experiment. Recall that $\delta=0$ in (27) corresponds to the temperature-independent $\omega_{j}(\bm{k})$, while for $\delta=1$, $\omega_{j}(\bm{k},T)$ is strongly dependent on $T$ according to (25).
[…]
@@ V Discussion and Summary
We generalized both approaches to the multi-sublattice collinear ferrimagnetic. In the MFA framework, we accounted for two magnetic sub-lattices and derived coupled equations Eq. (15) for the $T$- dependence of the mean spins for each sublattice, and the analytical expression in Eq. (16a) for the Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF}}$.
Within SWA, we accounted for multiple magnon frequency spectra branches and derived Eq. (20b), which describes the suppression of $\overline{S}^{z}(T)$ by spin-wave excitations. Given that the exact derivation of the temperature dependence of the magnon frequencies via the diagrammatic technique is too cumbersome for the level of the approximations that we used in this work, we suggested a phenomenological description of $\omega_{j}(\bm{k},T)$, Eq. (27), that employs one tuning parameter $\delta$.
The central theoretical part of our paper focuses on the unifying SW and MF approximations for the multi-sublattice ferrimagnets. This approach culminates in Eq. (30) which connects $\overline{S}^{z}(T)$ with the total amount of magnons in the ferromagnetic and antiferromagnetic branches. Formally speaking, Eq. (30) is a relatively simple generalization of unified SW and MFA Eq. (28a) for the simple one-sublattice ferromagnetics such as EuO or EuS analyzed in Ref. [6]. In Eq. (30), we replaced the spin of a single ion $S$ in EuO with the total spin $\overline{S}^{z}(0)$ in the elementary cell of a multi-sublattice ferrimagnetic. Additionally, we substituted the total number of magnons $\mathcal{N}$ in a single FM branch with the value $\mathcal{N}_{+}-\mathcal{N}_{-}$, which accounts for magnons from all FM and AFM branches in the multi-sublattice ferrimagnets. These replacements are not justified by some consistent procedure that is expected to lead to much more complicated basic equations. A price for simplicity of the current version of a theory for magnetization in multi-sublattice ferrimagnets is the tuning parameter $\delta$ in Eq. (27) for $\omega_{j}(\bm{k},T)$, which was chosen to require the theoretical Curie temperature to match its experimental value.
The unified SW-MFA minimal model with the chosen tuning parameter $\delta=0.428$ demonstrate good quantitative agreement between the theoretical $\overline{S}^{z}(T)$ and experimental magnetization temperature dependencies of YIG in the entire temperature range from $T=0$ K to $T=T^{\rm Exp}_{\rm{}_{C}}$.
@@
We thus definitely demonstrated that $M(T)$ scales in YIG nearly normally as $\sqrt{T_{\rm{}_{C}}-T}$ for all temperatures. This behavior contrasts sharply with the anomalous scaling of $M(T)$ in simple ferromagnets EuO and EuS, where $\beta\approx 1/3$[6]. This fact highlights the need for further investigation of the magnetization of different materials.
Our theoretical approach marks a significant advancement in the description of magnetic systems. We demonstrated the effectiveness of unifying two simple theoretical approximations for the description of complex magnetic materials. We believe that the applicability of our SW-MFA minimal model is not limited to YIG but may be extended to various multi-sublattice ferrimagnets and antiferromagnets. A further generalization of our minimal model will allow us to study more complex magnetic structures, such as Gadolinium Gallium Garnet (GGG, Gd3Ga5O12) at low temperatures [38, 39, 40, 41]. GGG has the same crystal structure and lattice parameters as YIG, making it an ideal substrate material for the epitaxial films of YIG. These layers have recently been used in research on magnon lifetime and transport at millikelvin temperatures [42, 16, 43, 44], paving the way for advancements in quantum magnonics.